The area of a triangle is 130 square feet. If the height is 14 less than the base, find the length of the height and base of the triangle.
height= ??
Base=??

Answer by stanbon(75887)

The area of a triangle is 130 square feet. If the height is 14 less than the base, find the length of the height and base of the triangle.
Formula:: Area = (1/2)(base)(height)
-----------------------
130 = (1/2)(x)(x-14)
-----
x^2 - 14x - 260 = 0
----
Ans: x = 24.58.... (base)
x-14 = 10.58.. (height)
===========================
Cheers,
Stan H.
------------------